# TPP-SD: Accelerating Transformer Point Process Sampling with Speculative Decoding

### v2

@@ Abstract
We propose TPP-SD, a novel approach that accelerates Transformer temporal point process (TPP) sampling by adapting speculative decoding (SD) techniques from language models. By identifying the structural similarities between thinning algorithms for TPPs and speculative decoding for language models, we develop an efficient sampling framework that leverages a smaller draft model to generate multiple candidate events, which are then verified by the larger target model in parallel. TPP-SD maintains the same output distribution as autoregressive sampling while achieving significant acceleration. Experiments on both synthetic and real datasets demonstrate that our approach produces samples from identical distributions as standard methods, but with 2-6$\times$ speedup. Our ablation studies analyze the impact of hyperparameters such as draft length and draft model size on sampling efficiency. TPP-SD bridges the gap between powerful Transformer TPP models and the practical need for rapid sequence sampling.
@@ 1 Introduction
Temporal point processes (TPPs) (daley2003introduction, ) are essential stochastic models for describing discrete event occurrence patterns in continuous time. Classical models include the Poisson process (kingman1992poisson, ), which is history-independent, and the Hawkes process (hawkes1971spectra, ), which incorporates historical dependencies. Recently, advancements in deep learning have facilitated the integration of deep models with point processes, significantly enhancing their expressive capabilities in capturing history dependence. Notable examples include models based on traditional RNNs (du2016recurrent, ), LSTMs (mei2017neural, ), and Transformers (simiao2020transformer, ; zhang2020self, ). RNN-type methods often suffer from vanishing or exploding gradients (pascanu2013difficulty, ), limiting their performance. In contrast, Transformer TPPs have gained popularity for their ability to capture long-term dependencies and support parallel training.
@@
[…]
@@ 3 Related Works
In this section, we discuss related works on point process sampling. In traditional statistical point processes, the CIF parameterization is generally preferred because the CDF is not a convenient way to specify history-dependent point processes. In contrast, the CIF provides a more convenient way to specify how the present depends on the past. Therefore, sampling methods in the statistical point process field are generally based on the CIF. Inverse method (rasmussen2018lecturenotestemporalpoint, ) simulates a unit-intensity Poisson process and transforms it into the desired point process using the inverse of the integrated CIF, leveraging the time-rescaling theorem (brown2002time, ). However, this method is not suitable for complex processes that do not have analytical forms for the inverse of integrated CIF. Another more easily implementable sampling method is the thinning algorithm, which was originally proposed for sampling from inhomogeneous Poisson processes (lewis1979simulation, ), and later adapted for history-dependent processes (ogata1981lewis, ). With the development of the deep neural point process, an increasing number of works have started using the CDF parameterization. For example, (shchur2019intensity, ) used normalizing flows and mixtures of log-normal densities to model the CDF of the inter-event interval. Although this approach was initially proposed for training convenience, another advantage it brings is facilitating sampling. This is because we can easily sample the next event directly from the CDF (panosdecomposable, ). Whether based on CIF or CDF, both of the above methods for sampling history-dependent point processes can only be performed sequentially, leading to inefficient sampling. The key difference of this work is the introduction of parallelization in sampling, which accelerates the sampling process.
@@ 4.2 CDF-based Transformer TPPs
Our proposed CDF-based Transformer TPP model is illustrated in Fig.˜1(a). This model can be formally represented as =$\mathcal{M}=\{\mathcal{E},g(\tau|\cdot),f(k|\cdot)\}$, where $\mathcal{E}$ corresponds to the encoder and $g(\tau|\cdot),f(k|\cdot)$ correspond to the decoder. The model is trained by maximizing the log-likelihood in Eq.˜2.
@@ 4.3 TPP-SD
[…]
@@ 5.1 Evaluation of Sampling Quality and Speed
[…]
@@
[…]
@@ 5.2 Experimental Results on Synthetic Data
We consider three synthetic datasets: inhomogeneous Poisson, univariate Hawkes, and multivariate Hawkes processes, each with 1000 sequences within the time window [0,100][0,100][ 0 , 100 ]. For each dataset, we train an 8-head, 20-layer target model and a 1-head, 1-layer draft model. We compare two sampling approaches: AR sampling using only the target model versus TPP-SD which combines both target and draft models as elaborated in Section˜4.3. Details on data simulation procedures, data splitting, and experimental settings are provided in Sections˜B.1 and C.3.1.
@@ Results.
[…]
In addition, architectural variations are evident: AttNHP achieves the highest acceleration ratios despite having the slowest wall times; SAHP shows the lowest speedup but maintains the fastest wall times; and THP strikes a balance between acceleration ratio and runtime. These results suggest that the underlying architecture plays a significant role in determining the effectiveness of acceleration. A detailed discussion of these differences is provided in Section˜D.2.
@@
[…]
@@ 5.4 Ablation Studies
[…]
@@
We investigate the impact of draft model size on performance by fixing the target model to an 8-head, 20-layer Transformer and evaluating three draft configurations: 1-head-1-layer, 2-head-4-layer, and 4-head-6-layer. Experiments are conducted on the Multi-Hawkes and Taobao datasets. Results show that increasing draft model size preserves sampling quality—as measured by $\Delta\mathcal{L}$ and $D$—and improves the acceptance rate $\alpha$, but reduces the speedup ratio $S_{\text{AR/SD}}$. Notably, the 1-head-1-layer configuration achieves the highest speedup without compromising sample quality.
@@ 6 Conclusions
[…]

### v3

@@ Abstract
[…]
@@ 1 Introduction
Temporal point processes (TPPs) (daley2003introduction, ) are essential stochastic models for describing discrete event occurrence patterns in continuous time. Classical models include the Poisson process (kingman1992poisson, ), which is history-independent, and the Hawkes process (hawkes1971spectra, ; zhou2020auxiliary, ; zhou2022efficient, ), which incorporates historical dependencies. Recently, advancements in deep learning have facilitated the integration of deep models with point processes, significantly enhancing their expressive capabilities in capturing history dependence (meng2024transfeat, ). Notable examples include models based on traditional RNNs (du2016recurrent, ), LSTMs (mei2017neural, ), and Transformers (simiao2020transformer, ; zhang2020self, ; meng2024interpretable, ). RNN-type methods often suffer from vanishing or exploding gradients (pascanu2013difficulty, ), limiting their performance. In contrast, Transformer TPPs have gained popularity for their ability to capture long-term dependencies and support parallel training.
@@
Specifically, we make the following key contributions in this work: (1) We identify a strong similarity between the thinning algorithm used in point process sampling and the SD technique in LLMs. (2) We propose TPP-SD, which successfully applies SD to Transformer TPPs, resulting in accelerated sampling. (3) Extensive experiments are conducted on both synthetic and real datasets. The results show that TPP-SD and standard autoregressive sampling produce samples from the same distribution, but TPP-SD achieves a speedup of approximately $2\text{--}6\times$. Ablation studies analyze the impact of different hyperparameters.
@@ 3 Related Works
In this section, we discuss related works on point process sampling. In traditional statistical point processes, the CIF parameterization is generally preferred because the CDF is not a convenient way to specify history-dependent point processes. In contrast, the CIF provides a more convenient way to specify how the present depends on the past. Therefore, sampling methods in the statistical point process field are generally based on the CIF. Inverse method (rasmussen2018lecturenotestemporalpoint, ) simulates a unit-intensity Poisson process and transforms it into the desired point process using the inverse of the integrated CIF, leveraging the time-rescaling theorem (brown2002time, ), and (shchur2020fast, ) applies the inverse map in parallel to accelerate this process. However, this method is not suitable for complex processes that do not have analytical forms for the inverse of integrated CIF. Another more easily implementable sampling method is the thinning algorithm, which was originally proposed for sampling from inhomogeneous Poisson processes (lewis1979simulation, ), and later adapted for history-dependent processes (ogata1981lewis, ; Wang_Cheng_Yuan_Xu_2023, ). With the development of the deep neural point process, an increasing number of works have started using the CDF parameterization. For example, (shchur2019intensity, ) used normalizing flows and mixtures of log-normal densities to model the CDF of the inter-event interval. Although this approach was initially proposed for training convenience, another advantage it brings is facilitating sampling. This is because we can easily sample the next event directly from the CDF (panosdecomposable, ). Whether based on CIF or CDF, both of the above methods for sampling history-dependent point processes can only be performed sequentially, leading to inefficient sampling. The key difference of this work is the introduction of parallelization in sampling, which accelerates the sampling process.
@@ 4.2 CDF-based Transformer TPPs
Our proposed CDF-based Transformer TPP model is illustrated in Fig.˜1(a). This model can be formally represented as $\mathcal{M}=\{\mathcal{E},g(\tau|\cdot),f(k|\cdot)\}$, where $\mathcal{E}$ corresponds to the encoder and $g(\tau|\cdot),f(k|\cdot)$ correspond to the decoder. The model is trained by maximizing the log-likelihood in Eq.˜2.
@@ 4.3 TPP-SD
where $\text{norm}(\cdot)$ denotes the normalization. The correctness and necessity of sampling from the adjusted distribution have been established in the original speculative decoding work for LLMs (leviathan2023fast, ). In this work, we extend the proof to the TPP domain, as presented in Section˜A.2.
Eq.˜4 is a discrete distribution, and its normalization is relatively easy to implement, but Eq.˜3 is a continuous distribution, and normalizing it is much more difficult because we need to compute the normalizing constant $\int\max(0,g_{T}(\tau_{i+l}|\cdot)-g_{D}(\tau_{i+l}|\cdot))d\tau_{i+l}$. This is also the main difference between TPP-SD and LLM-SD, as continuous distributions are not involved in the LLM application. Inspired by (wang2024continuousspeculativedecodingautoregressive, ), we employ an acceptance-rejection sampling scheme (casella2004generalized, ) to sample from $g^{\prime}(\tau_{i+l}|\cdot)$, as outlined in Theorem˜1 (refer to Section˜A.3 for the proof).
@@ 5.1 Evaluation of Sampling Quality and Speed
For synthetic data where the ground truth is known, the time-rescaling theorem (brown2002time, ) states that for event times $\{t_{i}\}_{i=1}^{n}$ generated by a point process with CIF $\lambda^{*}(t)$, if the model is correctly specified, the transformed inter-event intervals $z_{i}=\int_{t_{i-1}}^{t_{i}}\lambda^{*}(\tau)\,d\tau$ are i.i.d. samples from $\text{Exponential}(1)$. Therefore, we apply the ground-truth CIF to transform the samples generated by AR sampling and TPP-SD, and then compute the KS statistic $D_{\text{KS}}$ to assess the conformity of generated samples to the ground-truth distribution. A smaller $D_{\text{KS}}$ indicates higher sample quality. The detailed computation of $D_{\text{KS}}$ is provided in Section˜A.4.
@@
For real data where the ground truth is unknown, the KS statistic cannot be computed. Instead, we assess the sampling consistency between AR sampling and TPP-SD using the Wasserstein distance $D_{\text{WS}}$. Specifically, using the first $M$ events as history, we perform $N$ independent repetitions of sampling $(M+1)$-th event, yielding $\{(t_{i}^{\text{AR}},k_{i}^{\text{AR}})\}_{i=1}^{N}$ from AR sampling and $\{(t_{i}^{\text{SD}},k_{i}^{\text{SD}})\}_{i=1}^{N}$ from TPP-SD. For the temporal distribution, we compute $D_{\text{WS}}^{t}$, the 1-Wasserstein distance between the empirical distributions of $\{t_{i}^{\text{AR}}\}_{i=1}^{N}$ and $\{t_{i}^{\text{SD}}\}_{i=1}^{N}$, using ot.wasserstein_1d from the POT library (flamary2021pot, ). For the event type distribution, we compute $D_{\text{WS}}^{k}$, the earth mover’s distance between the empirical distributions of $\{k_{i}^{\text{AR}}\}_{i=1}^{N}$ and $\{k_{i}^{\text{SD}}\}_{i=1}^{N}$, using ot.emd2 from the same library. A smaller $D_{\text{WS}}$ indicates higher sample quality.
@@ 5.2 Experimental Results on Synthetic Data
We consider three synthetic datasets: inhomogeneous Poisson, univariate Hawkes, and multivariate Hawkes processes, each with 1000 sequences within the time window $[0,100]$. For each dataset, we train an 8-head, 20-layer target model and a 1-head, 1-layer draft model. We compare two sampling approaches: AR sampling using only the target model versus TPP-SD which combines both target and draft models as elaborated in Section˜4.3. Details on data simulation procedures, data splitting, and experimental settings are provided in Sections˜B.1 and C.3.1.
@@ Results.
TPP-SD consistently achieves faster execution times and $1.3\text{--}5.7\times$ speedup across all datasets and encoder architectures. These results demonstrate that TPP-SD maintains the same high sampling fidelity as AR sampling while providing significant improvements in sampling efficiency.
In addition, architectural variations are evident: AttNHP achieves the highest acceleration ratios despite having the slowest wall times; SAHP shows the lowest speedup but maintains the fastest wall times; and THP strikes a balance between acceleration ratio and runtime. These results suggest that the underlying architecture plays a significant role in determining the effectiveness of acceleration. A detailed discussion of these differences is provided in Section˜D.4.
@@
Table˜2 shows TPP-SD consistently delivers high-fidelity sampling across all real datasets and encoder architectures. Similar to the observations on synthetic data, $\Delta\mathcal{L}^{\text{real}}$ remains low, and both $D_{\text{WS}}^{t}$ and $D_{\text{WS}}^{k}$ approach zero, indicating strong alignment in both temporal and type distributions. Moreover, TPP-SD achieves $1.6\text{--}5.9\times$ faster execution across all settings. These results confirm that TPP-SD matches the sampling fidelity of AR sampling while offering substantial gains in efficiency.
@@ 5.4 Ablation Studies
To identify the optimal draft length $\gamma$ for TPP-SD, we evaluate performance across datasets within the time window $[0,100]$ for $\gamma\in[1,60]$. Fig.˜3 shows that variations in $\gamma$ have negligible impacts on the metrics $\Delta\mathcal{L}$ and $D$, confirming the distributional equivalence of TPP-SD and AR sampling. However, acceptance rate $\alpha$ decreases as $\gamma$ increases, and speedup $S_{\text{AR/SD}}$ peaks before declining—falling below $1\times$ for excessively large draft lengths due to the overhead of evaluating rejected drafts. Moderate draft lengths (e.g., $\gamma\approx 5\text{--}15$) yield the highest speedup across all datasets.
@@
We investigate the impact of draft model size on performance by fixing the target model to an 8-head, 20-layer Transformer and evaluating three draft configurations: 1-head-1-layer, 2-head-4-layer, and 4-head-6-layer. Experiments are conducted on the Multi-Hawkes and Taobao datasets. Results show that increasing draft model size preserves sampling quality—as measured by $\Delta\mathcal{L}$ and DD—and improves the acceptance rate $\alpha$, but reduces the speedup ratio $S_{\text{AR/SD}}$. Notably, the 1-head-1-layer configuration achieves the highest speedup without compromising sample quality.
@@ 6 Conclusions
In this work, we introduce the original SD framework from the LLM domain into the context of TPP sampling. By identifying structural similarities between the thinning algorithm in TPPs and speculative decoding in LLMs, we develop an efficient framework that employs a lightweight draft model to propose candidate events for verification by the target model. TPP-SD significantly improves sampling efficiency by $2\text{--}6\times$ while preserving distributional consistency with AR sampling, thus bridging the gap between the expressive power of Transformer TPPs and the need for efficient sequence generation in practical applications. As future work, we plan to further optimize TPP-SD by incorporating advancements in SD. Specifically, we aim to integrate the draft mechanism into the target model (10.5555/3692070.3692273, ), or perform speculative decoding at the feature level rather than the event level (10.5555/3692070.3693232, ).
Acknowledgments
This work was supported by the NSFC Project (No.62576346), the MOE Project of Key Research Institute of Humanities and Social Sciences (22JJD110001), the fundamental research funds for the central universities, and the research funds of Renmin University of China (24XNKJ13), and Beijing Advanced Innovation Center for Future Blockchain and Privacy Computing.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope. See the Abstract and Introduction sections.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitation of the work is discussed in the appendix.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: For all theoretical results, the paper provides the corresponding proofs in the appendix.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper fully discloses all the information needed to reproduce the main experimental results in the paper. See the Experiments section.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide the data and code in the supplemental material to reproduce the main experimental results.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We specify all the training and test details necessary to understand the results. See the Experiments section.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report the statistical significance of the experiments. In the figures presenting the experimental results, we provide the standard deviations across multiple runs.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: All experiments are conducted on the server whose details are provided in the paper. See the Experiments section.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper conform with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [No]
  49. Justification: This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All code, models, and datasets mentioned in the text are appropriately cited with their original papers.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: New assets introduced in the paper, such as code, are well documented. The documentation is provided alongside the assets in the supplementary material.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.